OpenStudy (anonymous):

Find y" for y = sin2(x)

OpenStudy (anonymous):

@satellite73 can u help?

OpenStudy (wikiemol):

I am assuming you mean \[y = \sin^2(x)\] in which case use the product rule. which is d/dx f(x)g(x) = f'(x)g(x) + f(x)g'(x)

OpenStudy (anonymous):

i got this is it correct? y = 2(cos^2 x - sin^2 x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

cos^2(x)-sin^2(x)=cos2(x) y'=-2sin2x y"=-4cos2x

OpenStudy (anonymous):

ok